# Electromagnetic characteristics as probes into the inner structures of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states

### v1

@@ II.2.3 Coupled-channel analysis
As in the case of SSDD wave mixing analysis, a rigorous study of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks under coupled-channel effects requires the spatial wave functions of all coupled channels as essential inputs. In the subsequent analysis, we continue to employ three representative binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV to systematically evaluate the corresponding magnetic moments. The results for the magnetic moments of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates in coupled-channel analysis are summarized in Fig. 4.
A systematic comparison of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from single-channel and coupled-channel analyses reveals that coupled-channel effects can alter their numerical values. In particular, while the magnetic moments of the $\Xi_{c}D_{s}^{*}$ state with $J^{P}=1/2^{-}$ are identical for $I_{z}=+1/2$ and $I_{z}=-1/2$ in single-channel analysis, they become distinct once coupled-channel effects are included.
@@ III The M1 radiative decay behavior of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecules
Building upon our investigation of the magnetic moments of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks, we now turn to their M1 radiative decay properties. As a fundamental electromagnetic observable, the M1 radiative decays probe the transition amplitudes between the hadronic states and offer critical insight into the inner wave function configurations. In this section, we systematically analyze the M1 radiative decay behavior of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states within the constituent quark model using three complementary approaches: single-channel analysis, SSDD wave mixing analysis, and coupled-channel analysis.
@@
In the following, we outline the formalism for calculating the M1 radiative decay width between the hadronic molecules within the constituent quark model. The M1 radiative decay width between the hadronic molecules for the ${[AB]}\rightarrow{[AB]}^{\prime}\gamma$ process, denoted as $\Gamma_{{[AB]}\rightarrow{[AB]}^{\prime}\gamma}$, is given by the expression Wang:2016dzu ; Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Wang:2023bek ; Lai:2024jfe ; Zhang:2025ame :
@@
As indicated in Eq. (3.6), the central quantity in the calculation of the M1 radiative decay width between the hadronic states is the corresponding transition magnetic moment, denoted as $\mu_{{[AB]}\to{[AB]}^{\prime}\gamma}$. In this work, we calculate the transition magnetic moments between the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states using the following expression Wang:2016dzu ; Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Wang:2023bek ; Lai:2024jfe ; Zhang:2025ame :
@@
where $Y_{lm}(\Omega)$ is the spherical harmonic function, $L_{n}^{l+\frac{1}{2}}(x)$ is the associated Laguerre polynomial, and $n$, $l$, and $m$ are the radial, orbital, and magnetic quantum numbers of the discussed hadrons, respectively. By fitting their experimental mass spectra ParticleDataGroup:2024cfk , we determine the oscillator parameters as Luo:2023sne :
@@
We now investigate the M1 radiative decay behavior of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states by including contributions from the $D$-wave channels. A comprehensive set of the corresponding decay widths of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates obtained in SSDD wave mixing analysis is provided in Table 2.
Compared to single-channel results, the inclusion of SSDD wave mixing effects has a negligible impact on the M1 radiative decay widths of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates. This result aligns with our earlier analysis of their magnetic moments and reflects very small $D$-wave admixture in the total spatial wave function, which in turn leads to only minor changes in the electromagnetic observables.

### v2

@@ II.2.3 Coupled-channel analysis
As in the case of $S$-$D$ wave mixing analysis, a rigorous study of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks under coupled-channel effects requires the spatial wave functions of all coupled channels as essential inputs. In the subsequent analysis, we continue to employ three representative binding energies of 0.5-0.5, 6.0-6.0, and 12.0-12.0 MeV to systematically evaluate the corresponding magnetic moments. The results for the magnetic moments of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates in coupled-channel analysis are summarized in Fig. 4.
A systematic comparison of the magnetic moments of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks obtained from single-channel and coupled-channel analyses reveals that coupled-channel effects can alter their numerical values. In particular, while the magnetic moments of the $\Xi_{c}D_{s}^{*}$ state with $J^{P}={1/2}^{-}$ are identical for $I_{z}=+1/2$ and $I_{z}=-1/2$ in single-channel analysis, they become distinct once coupled-channel effects are included. This arises primarily because some channels within the coupled system may exhibit different magnetic moments depending on $I_{z}$. In addition, the stability of the isospin-dependent effects in the electromagnetic observables of the molecular states observed in the coupled-channel calculations is governed by two key factors: (i) the probability percentages of each channel within the coupled system, which are determined by the mass spectrum of the coupled system, and (ii) the magnetic moments and the transition magnetic moments of the individual channels, which depend on the flavor and spin structures of the constituent hadrons.
@@ III The M1 radiative decay behavior of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecules
Building upon our investigation of the magnetic moments of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks, we now turn to their M1 radiative decay properties. As a fundamental electromagnetic observable, the M1 radiative decays probe the transition amplitudes between the hadronic states and offer critical insight into the inner wave function configurations. In this section, we systematically analyze the M1 radiative decay behavior of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states within the constituent quark model using three complementary approaches: single-channel analysis, $S$-$D$ wave mixing analysis, and coupled-channel analysis.
@@
In the following, we outline the formalism for calculating the M1 radiative decay width between the hadronic molecules within the constituent quark model. The M1 radiative decay width between the hadronic molecules for the ${[AB]}\rightarrow{[AB]}^{\prime}\gamma$ process, denoted as $\Gamma_{{[AB]}\rightarrow{[AB]}^{\prime}\gamma}$, is given by the expression Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Wang:2023bek ; Lai:2024jfe ; Zhang:2025ame ; Dey:1994qi ; Majethiya:2009vx ; Majethiya:2011ry :
@@
As indicated in Eq. (3.6), the central quantity in the calculation of the M1 radiative decay width between the hadronic states is the corresponding transition magnetic moment, denoted as $\mu_{{[AB]}\to{[AB]}^{\prime}\gamma}$. In this work, we calculate the transition magnetic moment between the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states using the following expression Li:2021ryu ; Zhou:2022gra ; Wang:2022tib ; Wang:2023bek ; Lai:2024jfe ; Zhang:2025ame :
@@
where $Y_{lm}(\Omega)$ is the spherical harmonic function, $L_{n}^{l+\frac{1}{2}}(x)$ is the associated Laguerre polynomial, and $n$, $l$, and $m$ are the radial, orbital, and magnetic quantum numbers of the discussed hadrons, respectively. By fitting their experimental mass spectra ParticleDataGroup:2024cfk , we determine the oscillator parameters $\beta$ as Luo:2023sne :
@@
Here, it is important to note that the oscillator parameters $\beta$ characterize the spatial wave functions of the baryons $\Xi_{c}^{(^{\prime},*)}$ and the mesons $D^{(*)}_{s}$ serve as key inputs in evaluating the electromagnetic properties of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states. In this work, the $\beta$ values for each hadrons are determined by fitting their experimental mass spectra. Given the uncertainties associated with these parameters, it is necessary to assess their impact on the electromagnetic properties of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states. Following a similar approach to the treatment of the constituent quark masses, we examine two typical radiative decay processes by varying $\beta$ within $\pm 10\%$ in single-channel analysis. The numerical results are as follows:
$\Gamma_{[\Xi_{c}^{{}^{\prime}}D_{s}]|{1}/{2}^{-}\rangle\to[\Xi_{c}D_{s}]|{1}/{2}^{-}\rangle\gamma}^{I_{z}=1/2}$ $=$ $20.52^{+1.05}_{-4.08}{~\rm keV},$
$\Gamma_{[\Xi_{c}^{*}D_{s}]|{3}/{2}^{-}\rangle\to[\Xi_{c}D_{s}]|{1}/{2}^{-}\rangle\gamma}^{I_{z}=1/2}$ $=$ $70.61^{+5.74}_{-17.15}{~\rm keV}.$
In the calculations, the binding energies of both the initial and final molecular states are taken to be 6.0-6.0 MeV. The resulting uncertainties in the radiative decay widths arising from the variation of $\beta$ are indicated as the superscripts and subscripts attached to the central values. While the numerical values of the radiative decay widths may shift to some extent when the oscillator parameters are varied, the order of magnitude generally remains stable. This point has been discussed in more detail in our previous work Wang:2022nqs .
We now investigate the M1 radiative decay behavior of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states by including contributions from the $D$-wave channels. A comprehensive set of the corresponding decay widths of the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates obtained in $S$-$D$ wave mixing analysis is provided in Table 3.
Compared to single-channel results, the inclusion of $S$-$D$ wave mixing effects has a negligible impact on the M1 radiative decay widths of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular candidates. This result aligns with our earlier analysis of their magnetic moments and reflects very small $D$-wave admixture in the total spatial wave function, which in turn leads to only minor changes in the electromagnetic observables.
@@
To provide further guidance for experimental searches, we present the following discussion. Analogous to the well-established double-charm tetraquark state $T_{cc}(3875)^{+}$ LHCb:2021auc ; LHCb:2021vvq , we suggest that the predicted $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular pentaquarks can be produced in the $pp$ collisions at the LHCb. As argued in Refs. Ozdem:2025olj ; Ozdem:2026wmf , the hadrons with relatively large magnetic moments are expected to couple more strongly to external electromagnetic fields, thereby enhancing their production in the photon-induced channels. The magnetic moments of the order of 1 $\mu_{N}$ obtained for the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states indicate that the electromagnetic production mechanisms may be within experimental reach, for instance via $\gamma p\to P_{ccs\bar{s}}X$ at facilities such as GlueX or future electron-ion colliders. A more direct probe of the electromagnetic structures of the $P_{ccs\bar{s}}$ states lies in their radiative decays. These can be observed in processes such as $e^{+}e^{-}\to P_{ccs\bar{s}}\gamma X$ or $pp\to P_{ccs\bar{s}}\gamma X$, at experiments like Belle II and LHCb, with the radiative decay widths serving as the key observables. To enable a full reconstruction chain, the decays of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states can be traced through the sequential decays of their constituents. The constituent baryons may be reconstructed using the main decay channels, for example, $\Xi_{c}^{*}\to\Xi_{c}\pi$, $\Xi_{c}^{\prime}\to\Xi_{c}\gamma$, $\Xi_{c}\to\Xi\pi$, followed by $\Xi\to\Lambda\pi$ and $\Lambda\to p\pi$. The constituent mesons can be identified via decays such as $D_{s}^{*}\to D_{s}\gamma\to\phi(1020)\pi\gamma\to KK\pi\gamma$.
@@
In this work, we focus specifically on the electromagnetic properties of the $\Xi_{c}^{(^{\prime},*)}D^{(*)}_{s}$ molecular states. Our results demonstrate that these observables are highly sensitive to the spin-parity assignments of the hadronic molecules, thus offering valuable theoretical guidance for future experimental determination of their spin-parity quantum numbers. It would be particularly valuable to investigate the electromagnetic properties of the compact double-charm hidden-strangeness pentaquarks in future studies, as such work could help clarify the inner configurations of the double-charm hidden-strangeness pentaquarks. Currently, the experimental study of the electromagnetic properties of the hadrons, particularly for the exotic states, faces significant challenges. Nevertheless, recent progress in lattice QCD calculations has begun to address the electromagnetic properties of the exotic states Vujmilovic:2025czt , providing valuable theoretical benchmarks and motivating further experimental and theoretical investigations. It is worth noting that the strong decay widths of these molecular states might be significantly larger than their radiative decay widths. Looking ahead, we strongly encourage further studies employing a variety of models and methods to explore other properties of these hadronic molecules, including their strong decay behaviors. A comprehensive understanding of their inner structures and eventual experimental identification will ultimately rely on the combination of different types of observables, each offering complementary insights.